Low level commissioning 2005 2006 collaboration meeting baton rouge mark krasberg april 12 th 2006 l.jpg
This presentation is the property of its rightful owner.
Sponsored Links
1 / 28

Low Level Commissioning 2005/2006 Collaboration Meeting, Baton Rouge Mark Krasberg April 12 th , 2006 PowerPoint PPT Presentation


  • 171 Views
  • Updated On :
  • Presentation posted in: Pets / Animals

Low Level Commissioning 2005/2006 Collaboration Meeting, Baton Rouge Mark Krasberg April 12 th , 2006. Commissioning a DOM at The Testing DAQ Level. Commissioning plan pretty much followed. Could not study string 29 during freeze-in

Related searches for Low Level Commissioning 2005/2006 Collaboration Meeting, Baton Rouge Mark Krasberg April 12 th , 2006

Download Presentation

Low Level Commissioning 2005/2006 Collaboration Meeting, Baton Rouge Mark Krasberg April 12 th , 2006

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Low level commissioning 2005 2006 collaboration meeting baton rouge mark krasberg april 12 th 2006 l.jpg

Low Level Commissioning 2005/2006Collaboration Meeting,Baton RougeMark KrasbergApril 12th, 2006

Mark Krasberg

U of Wisconsin


Commissioning a dom at the testing daq level l.jpg

Commissioning a DOMat The Testing DAQ Level

Mark Krasberg

U of Wisconsin


Commissioning plan pretty much followed l.jpg

Commissioning plan pretty much followed

  • Could not study string 29 during freeze-in

  • Ended up taking more data than original plan (not a bad thing)

  • Cold temperature problem meant that not all elements of the plan could be followed

Mark Krasberg

U of Wisconsin


Problem quads l.jpg

Problem Quads

  • Leaking Quads

    • 12/25 String 29 all quads

    • 01/04 String 39 none (quad 7 SJB problem)

    • 01/08 String 38 leak-checked late (oscillations)

    • 01/14 String 30 quads 4, 7 (quad 14 SJB prob)

    • 01/18 String 40 quads 13 and 14

    • 01/22 String 50 leak-checked late (damaged)

    • 01/25 String 59 leak-checked late (swamped)

    • 01/29 String 49 quads 6, 9 and 10

  • Anomolies

    • Any DOM which exhibited anomolous behavior was unplugged

      • Several wire pairs had high currents and their quads were unplugged and periodically tested

  • Communications Issues

    • Encountered a cold temperature (cold reboot) problem

    • Issue was a moving target as DOMs cooled and changed temperatures

    • Problem seems to be understood (final fix not applied at Pole yet)

Mark Krasberg

U of Wisconsin


String 39 two week freeze in movie l.jpg

String 39 two-week freeze-in movie

Mark Krasberg

U of Wisconsin


Triboluminescence l.jpg

Triboluminescence

  • Had a longstanding plan on how to cope with the triboluminescence: (>40 kHz individual DOM rates!)

    • Use LC up.AND.down (“Triple Local Coincidence”)

    • Use Scaler Deadtime (reduces LC trigger rate)

  • The plan worked!

    • able to take over 1 terrabyte of data during season and hand-carry north

Mark Krasberg

U of Wisconsin


Slide7 l.jpg

Data-taking during Triboluminescence

Approximate maximum trigger rate for normal TestDAQ data-taking

frozen-in LC rate

Mark Krasberg

U of Wisconsin


Surface daq hardware l.jpg

Surface DAQ Hardware

  • Jan 15: string 21 hub stopped working, probably overheated

    • Did not operate all hubs simultaneously after this event until the power distribution and cooling fan problems in the TCH were solved

  • Jan 30: string 59 DSB card didn’t work – was replaced

  • String 59 DOMHub suffers from domhub-app JVM crashes. This is the only hub which has had more than one JVM crash with TestDAQ. RealDAQ has had problems with other hubs.

Mark Krasberg

U of Wisconsin


Cold temperature problems l.jpg

Cold Temperature Problems

  • DOMs found to have temperature-related problems (“cold reboot problem”)

  • Problems occur primarily at top of string (colder)

  • THIS MADE IT IMPOSSIBLE TO COMMISSION ANY DOMs at Pole

    • communications problems were a moving target

    • the list of DOMs which would not communicate properly would change daily, as the DOMs cooled down.

  • Particularly problemmatic temperature is ~-28C

Mark Krasberg

U of Wisconsin


Study of cold reboot problem in ice l.jpg

Study of cold reboot problem in ice

  • Used internal heating of mainboard to probe a temperature range (like in FATs)

  • String 29 “Farsta”

    • 57 reboots between -34.44 C and -27.25 C

    • 49 successful cold reboots

    • All reboots (8) failed between -31.25 C and -30.56 C

    • No successful reboots in this range

  • Just as during FAT testing, the problem is isolated to a specific, narrow temperature range

Mark Krasberg

U of Wisconsin


Effect of cold temperature problems on commissioning l.jpg

Effect of cold temperature problems on Commissioning

Mark Krasberg

U of Wisconsin


Able to deal with the expected unexpected problems fairly easily part of the plan l.jpg

Able to deal with the expected unexpected problems fairly easily(part of the plan!)

  • When we hook up DOMs in a new environment we often encounter problems

  • Each DOMHub had a full TestDAQ development workspace (ie all the code needed for data-taking could be modified and recompiled at pole).

  • Patches could be emailed through Iridium satellite

  • Patches could then be installed, recompiled, and tested within minutes

  • Feedback could be given to the north almost instantaneously using Iridium satellite – we did not have to wait for TDRS satellite. If we had, then the resulting code iterations would have gone on for days!

Mark Krasberg

U of Wisconsin


Communications now stable l.jpg

Communications now stable

  • TestDAQ softboot problem eliminated March 8.

    • DOM communications fix waiting to be deployed at pole

  • Data taking became stable with TestDAQ as of March 8th (600 DOMs at that time) except for occasional string59 domhub-app crashes.

  • DOMs can be commissioned since the communications problems have effectively been eliminated and data-taking is stable.

  • AMANDA mainboards have been included in the readout in almost all TestDAQ runs since February

Mark Krasberg

U of Wisconsin


High current doms l.jpg

High Current DOMs

  • Six high current wire pairs identified at pole

  • Four of the high currents mysteriously went away (seemingly not related to temperature or freeze-in). One took weeks to go away!

  • Two remaining high current pairs permanently plugged in at station close

    • DOMs on both of these pairs have suffered problems in the months since they were plugged in

    • In the future, high current pairs should not be used in normal data-taking for as long as is feasible (one year?)

Mark Krasberg

U of Wisconsin


High current dom pairs l.jpg

High Current DOM Pairs

  • High current pairs which are still high current pairs

  • | 39-07 | Hippophilia | 8b56bf63b908 | UP5P0888 | 0 | 3 |

  • | 39-08 | Shamal | 7b492c28a830 | TP5P0527 | 0 | 3 |

  • | 40-51 | Juneberry | ce9692f580c9 | UP4P0086 | 6 | 1 |

  • | 40-52 | Alfa_Romeo_Spider | 0a9dc8de00c3 | TP5Y0171 | 6 | 1 |

  • High current pairs which are no longer high current pairs:

  • | 49-01 | Ceres | 754dad1d3800 | UP4Y0030 | 0 | 0 |

  • | 49-02 | Marzahn | 60cad4124a82 | TP5Y0161 | 0 | 0 |

  • | 50-37 | Volkswagen_Beetle | 93c6fcb0fe68 | UP5Y0216 | 4 | 2 |

  • | 50-38 | Warthog | a095b31f89cf | TP4P0315 | 4 | 2 |

  • | 40-57 | Tortellini | d54044599535 | UP4P0140 | 7 | 0 |

  • | 40-58 | Cannelloni | 3dac398e0d9d | TP4P0111 | 7 | 0 |

  • these DOMs drew an extremely large current initially, before problem went away...

  • | 30-45 | Capybara | 0d42d3e12c9f | UP4P0276 | 5 | 2 |

  • | 30-46 | Asparagus | 44c3682eb25c | TP4P0217 | 5 | 2 |

  • The pair with 50-58 “Universitet” had a marginally high current at pole but was not flagged as high current – it’s LC was determined to be broken as soon as it was turned on.

Mark Krasberg

U of Wisconsin


Slide16 l.jpg

Mark Krasberg

U of Wisconsin


Slide17 l.jpg

40-51 “Juneberry” & 40-52 “Alfa_Romeo_Spider” poor comms

50-58 “Universitet” bad LC

39-08 “Shamal” bad LC

59-59 “Cosmology” poor flash

29-18 “Yurei” PMT?

38-59 “Blackberry” no HV

49-17 “Biometeorology” bad flash

30-23 “Peugeot_505” no power

50-36 “Ocelot” no power

29-59 “Auroraphobia” & 29-60 “Nix” no power

Mark Krasberg

U of Wisconsin


There are currently 16 strings not 9 l.jpg

There are currently 16 strings (not 9)!

hole: 21 29 30 38 39 40 49 50 59

strings: 1 1 2 1 2 2 2 3 2 total=16

we had 17 mini-strings in UW FAT #6!

broken LC connection between two DOMs

Problem DOMs

a very short string!

cannot be used for hard LC

Mark Krasberg

U of Wisconsin


Stf results for string 38 odd u doms l.jpg

STF results for string 38, odd (U) DOMs

Mark Krasberg

U of Wisconsin


38 59 blackberry failure l.jpg

38-59 “Blackberry” failure

Mark Krasberg

U of Wisconsin


38 59 blackberry failure zoom l.jpg

38-59 “Blackberry” failure zoom

Mark Krasberg

U of Wisconsin


Freeze in 49 55 fusilli dom was still in water on mar 22 l.jpg

Freeze-In: 49-55 “Fusilli”DOM was still in water on Mar 22

Mark Krasberg

U of Wisconsin


Slide23 l.jpg

String 49 – drill stopped for ~one hour at ~2300 meters

Mark Krasberg

U of Wisconsin


Summary of commissioned doms l.jpg

Summary of Commissioned DOMs

  • To date: the 300 DOMs for which STF ran successfully at Pole can be regarded as commissioned (low level)

    • when 49-55 “Fusilli” freezes-in STF will be run again, as per initial plan (don’t expect any problems)

  • Current Status: 600 out of 604 communicating on March 20

  • DOMs with serious problems

    • Three communicating DOMs not useful for hard LC data-taking (this affects 4th DOM)

    • 2 DOMs (40-51 “Juneberry” and 40-52 “Alfa_Romeo_Spider”) are high current, poorly communicating, have recently degraded and should be unplugged for special testing

  • DOMs with less serious problems

    • 2 DOMs (49-17 “Biometeorology” and 59-59 “Cosmology”) have flash problems

    • 29-18 “Yurei” suffered a 40% gain shift in March

    • 21-30 “Phenol” has sustained high dark rate problem, operating at lower gain

    • 21-13 “Wickueler” does not like to be turned off (discovered during this season)

  • Additional LC problem between 59-51 and 59-52 (broken connection during refreeze)

Mark Krasberg

U of Wisconsin


Slide25 l.jpg

Mark Krasberg

U of Wisconsin


Slide26 l.jpg

49-55 “Fusilli”

Mark Krasberg

U of Wisconsin


Slide27 l.jpg

Mark Krasberg

U of Wisconsin


Slide28 l.jpg

Mark Krasberg

U of Wisconsin


  • Login